What is the force between two protons that are separated by 1.0 mm? The charge of a proton is 1.6 × 10-19 C.


1
Expert's answer
2022-02-05T10:26:45-0500

"F=\frac{kq^2}{r^2}=2.3\cdot 10^{-22}~N."